# Non-commutative rings with infinitely many maximal subrings

### v1

@@ 2. Preliminaries
Let $T$ be a ring and $I$ and $J$ be left ideals of $T$. Then $I$ and $J$ are called similar left ideals if $T/I\cong T/J$ as left $T$-module. If $a\in T$ and define $(I:a):=\{x\in T\ |\ xa\in I\}$, then it is not hard to see that $I$ and $J$ are similar if and only if there exists $c\in T$ such that $I+Tc=T$ and $J=(I:c)$, see [13, Proposition 1.3.6]. Since for a left ideal $I$ of a ring $T$, we have the ring isomorphism $\mathbb{I}(I)/I\cong(End({}_{T}(T/I)))^{op}$, see [23, P. 14, 1.11], we immediately conclude that if $I$ and $J$ are similar left ideals of a ring $T$, then $\mathbb{I}(I)/I\cong\mathbb{I}(J)/J$ as rings, i.e., up to ring isomorphism, the ring $\mathbb{I}(I)/I$ is unique for similar left ideals; in particular $|\mathbb{I}(I)/I|=|\mathbb{I}(J)/J|$. In other words the cardinality of the ring $\mathbb{I}(I)/I$ is invariant up to similarity. Moreover, if $M$ is a maximal left ideal of a ring $T$, then we have the following facts too:
@@
    if $N$ is a (maximal) left ideal of $T$ which is similar to $M$, then $N$ is an ideal of $T$ if and only if $N=M$ if and only if $M$ is an ideal of $T$ .
To see the above facts, note that $(1)$ is an immediate consequences of [13, Proposition 1.3.6]. For $(2)$, let $M$ and $N$ be similar maximal left ideals of $T$ and assume that $N$ is an ideal of $T$. Since $T/N\cong T/M$ as left $T$-module we immediately conclude that $N=l.ann_{T}(T/N)=l.ann_{T}(T/M)\subseteq M$ and therefore $M=N$ (note that $N$ is a maximal left ideal of $T$, for $T/N\cong T/M$ as left $T$-module). Hence $(2)$ holds. Consequently, if $M$ is a maximal left ideal of $T$ which is not an ideal of $T$, and $[M]$ be the set of all (maximal) left ideals of $T$ which are similar to $M$, we deduce that $M]\subseteq Max_{l}(T)\setminus Max(T)$, by $(2)$, i.e., each (maximal) left ideals of $T$ which is similar to $M$ is not an ideal of $T$. Clearly, the relation similarity is an equivalence relation on the set of all left ideals of $T$ and in particular on $Max_{l}(T)$. By $(1)$ and $(2)$, it is obvious that if $M$ is a maximal left ideal of $T$, then $|[M]|=1$ (i.e., $M]=\{M\}$) if and only if $M$ is an ideal of $T$ (thus $M\in Max(T)\cap Max_{r}(T)$).
It is clear that whenever $I$ is a left ideal of a ring $T$ and $c\in T$, then $(I:c)=T$ if and only if $c\in I$. Moreover, $I\subseteq(I:c)$ if and only if $c\in\mathbb{I}(I)$. In particular, if $M$ is a maximal left ideal of a ring $T$, from these facts we immediately conclude that $(M:c)=M$ if and only if $c\in\mathbb{I}(M)\setminus M$. Hence $M]=\{M\}\cup\{(M:c)\ |\ c\in T\setminus\mathbb{I}(M)\}$. In the following main result we find a lower bound for $|[M]|$. Before present it, let us remind that from the previous observation for each $N\in[M]$, we have $|\mathbb{I}(M)/M|=|\mathbb{I}(N)/N|$. Now the following is in order.
@@
Let $T$ be a ring and $M\in Max_{l}(T)$ is not an ideal of $T$ (i.e., $M$ is not a maximal ideal of $T$). Then $|[M]|\geq|\mathbb{I}(M)/M|+1$. In particular, $|Max_{l}(T)\setminus Max(T)|\geq|\mathbb{I}(M)/M|+1$.
@@
Let $R=\mathbb{I}(M)$ and $D=R/M$. Since $M$ is not an ideal of $T$ we conclude that $R$ is a proper subring of $T$. Take a fixed element $x\in T\setminus R$. We claim that the function $c+M\longmapsto(M:x+c)$ is a one-one function from $D$ to $M]\setminus\{M\}$. Since $x\in T\setminus R$ and $c\in R$ we deduce that $x+c\in T\setminus R$ and therefore $(M:x+c)\in[M]\setminus\{M\}$, by the previous observation. Hence it remain to show that the function is one-one. Hence assume that $a,b\in R$ and $(M:x+a)=(M:x+b)$, we must show that $a+M=b+M$, i.e., $a-b\in M$. Let $t\in(M:x+a)=(M:x+b)$, then we conclude that $tx+ta,tx+tb\in M$. Therefore $t(a-b)\in M$, i.e., $(M:x+a)\subseteq(M:a-b)$. Now if $a-b\notin M$, then we conclude that $a-b\in R\setminus M$ and therefore (M:ab)=M(M:a-b)=M. Hence $(M:x+a)\subseteq M$ and therefore by maximality of $(M:x+a)$ we obtain that $(M:x+a)=M$ which is absurd, for $x+a\in T\setminus R$. Thus $a-b\in M$ and therefore we are done. ∎
As an immediate consequence of the previous theorem is that if $T$ is a ring and $Max_{l}(T)\setminus Max(T)\neq\emptyset$, then $|Max_{l}(T)\setminus Max(T)|\geq 3$, i.e., if a ring $T$ has a maximal left ideal which is not an ideal of $T$, then $T$ has at least three maximal left ideals which are not ideals of $T$.
@@
Let $T$ be a $K$-algebra over a field $K$. Then either $T$ is a left quasi duo ring or $T$ has at least $|K|+1$ maximal left ideals which are not ideals of $T$ (i.e., $|Max_{l}(T)\setminus Max(T)|\geq|K|+1$).
@@
Assume that $T$ is not a left quasi duo ring and $M$ be a maximal left ideal of $T$ which is not an ideal of $T$. Since $K$ is a field which is contained in the center of $T$, we conclude that $K\subseteq R:=\mathbb{I}(M)$ and clearly $K\cap M=0$. Thus $K$ embeds in $R/M$ and hence we are done by Theorem 2.1. ∎
@@
Let $T$ be an algebra over an infinite field $K$. Then either $T$ is a left quasi duo ring or $Max_{l}(T)\setminus Max(T)$ is infinite (in fact, $|Max_{l}(T)\setminus Max(T)|\geq|K|$).
@@
Thus, if $T$ is a ring and $Max_{l}(T)\setminus Max(T)$ is finite, then for each $M\in Max_{l}(T)\setminus Max(T)$, we deduce that $\mathbb{I}(M)/M$ is a finite field. Moreover, if in addition $T$ is a $K$-algebra over a field $K$ and $Max_{l}(T)\setminus Max(T)$ is finite, then $K$ is a finite field (note that $\mathbb{I}(M)/M$ contains a copy of $K$, for each maximal left ideal $M$ of $T$).
@@
By Theorem 2.1, note that if $Max_{l}(T)\setminus Max(T)=\bigcup_{i\in I}[M_{i}]$ be the partition of $Max_{l}(T)\setminus Max(T)$ in to similarity classes, then $|Max_{l}(T)\setminus Max(T)|\geq|I|+\sum_{i\in I}|\mathbb{I}(M)/M|$.
@@
It is clear that if $K$ is a field and $n>1$, then by the above results we have $|Max_{l}(\mathbb{M}_{n}(K))|\geq|K|+1$ and $|Max_{r}(\mathbb{M}_{n}(K))|\geq|K|+1$ (note, $\mathbb{M}_{n}(K)$ is a simple $K$-algebra). In particular, if $K$ is an infinite field then $\mathbb{M}_{n}(K)$ has at least $|K|$-many maximal left/right ideals (also see Remark 2.5). Next we want to use the notation and results of [27], to show that if $R$ is an infinite division ring and $n>1$, then $T=\mathbb{M}_{n}(R)$ has infinitely many maximal left/right ideals. To see this, note that if $M$ is a maximal left ideal of a ring $R$ and $T=\mathbb{M}_{n}(R)$, then for each $u\in R^{n}\setminus M^{n}$, the set $D(M,u):=\{X\in T\ |\ Xu\in M^{n}\}$ is a maximal left ideal of $T$ (in fact, $T/D(M,u)\cong(R/M)^{n}$ as left $T$-module). Moreover, $Max_{l}(T)=\{D(M,u)\ |\ M\in Max_{l}(R),\ u\in R^{n}\setminus M^{n}\}$, see [27, Theorem 1.2]. In particular, if $R$ is a division ring, then we deduce that $Max_{l}(\mathbb{M}_{n}(R))=\{D(0,u)\ |\ 0\neq u\in R^{n}\}$.
@@
Hence to show that for an infinite division ring $R$, $Max_{l}(\mathbb{M}_{n}(R))$ is infinite we must prove that there exist infinitely many $0\neq u\in R^{n}$ such that $D(0,u)$ are distinct maximal left ideals. By [27, Proposition 2.3], if $M$ is a maximal left ideal of a ring $R$, and $S=\mathbb{I}(M)$ be the idealizer of $M$ in $R$, and $u,v\in S^{n}$, then $D(M,u)=D(M,v)$ (in the ring $T=\mathbb{M}_{n}(R)$) if and only if there exists $c\in S\setminus M$ such that $v\equiv uc\ mod(M)$ (i.e. $v-uc\in M^{n}$). In particular, if $R$ is a division ring, since the only possible for $M$ is the zero ideal and in this case $S=R$, we deduce that for each $u,v\in R^{n}$, $D(0,u)=D(0,v)$ if and only if there exists $0\neq c\in R$ such that $v=uc$, (in other words $v$ and $u$ are right parallel vectors in $R^{n}$ as right $R$-module). Clearly, if $D$ is an infinite division ring and $n>1$, then $D^{n}$ has infinitely many right non-parallel vectors, thus we deduce the following fact.
@@
Let $R$ be an infinite division ring and $n>1$, then $T=\mathbb{M}_{n}(R)$ has infinitely many maximal left (resp. right) ideals.
@@
Let $R$ be a ring, $n>1$ and $T=\mathbb{M}_{n}(R)$. Then it is obvious that $Max_{l}(T)\cap Max(T)=\emptyset=Max_{r}(T)\cap Max(T)$. Moreover, the function $M\longmapsto M^{t}$ (where $M^{t}$ is the transpose of $M$, i.e., $M^{t}:=\{m^{t}\ |\ m\in M\}$), is a one-one corresponding from $Max_{l}(T)$ onto $Max_{r}(T)$. In particular, $|Max_{l}(T)|=|Max_{r}(T)|$.
@@
Let $T$ be a ring, then we have the following immediate facts:
@@
    $Max_{l}(T)\cap Max(T)=Max_{r}(T)\cap Max(T)=Prm_{l}(T)\cap Max_{l}(T)=Prm_{r}(T)\cap Max_{r}(T)=Prm_{l}(T)\cap Max_{r}(T)=Prm_{r}(T)\cap Max_{l}(T)$.
@@
    $Max(T)\setminus Max_{l}(T)=Max(T)\setminus Max_{r}(T)$.
Next let us remind the reader some immediate fact from maximal subrings and also some needed result from [1]. It is easy to see that a proper subring $S$ of a ring $T$ is maximal if and only if for each $x\in T\setminus S$, we have $S[x]=T$. Moreover, if $R$ is a proper subring of $T$ and there exists $\alpha\in T$ such that $R[\alpha]=T$, then by a natural use of Zorn’s Lemma, $T$ has a maximal subring $S$ with $R\subseteq S$ and $\alpha\notin S$. In particular, if $R\subsetneq T$ is a ring extension which is finite as ring (module) over $R$, then $T$ has a maximal subring which contains $R$. Now let $R$ be a ring and $\Delta:=\{(r,r)\ |\ r\in R\}$, be the diagonal subring of $T:=R\times R$. Then it is not hard too see that $\Delta[(1,0)]=T$. Therefore $T$ has a maximal subring which contains $\Delta$. In fact, subrings of $T$ which contains $\Delta$ has the form $\Delta(I):=\Delta+I\times I$, where $I$ is an ideal of $R$. In particular, $\Delta(I)$ is a maximal subring of $T$ if and only if $I$ is a maximal ideal of $R$, see [1, Theorem 2.7]. Therefore a ring $R$ is simple if and only if $\Delta$ is a maximal subring of $R\times R$. In the following we give some results from [1].
@@
    [1, Theorem 4.1]. Let $T$ be a ring and $A$ be a maximal right/left ideal of $T$ which is not an ideal of $T$. Then the idealizer of $A$ is a maximal subring of $T$. In particular, either a ring has a maximal subring or is a quasi duo ring.
@@
    [1, Proposition 4.2]. Let $R$ be a maximal subring of a ring $T$ which contains a maximal one-sided ideal $A$ of $T$ which is not an ideal of $T$. Then $R$ is the idealizer of $A$ in $T$.
@@
    [1, Theorem 4.4]. Let $T$ be a ring which is not a division ring. If $T$ is left/right primitive ring (in particular, if $T$ is a simple ring), then $T$ has a maximal subring.
@@
    [1, Theorem 3.1]. For each ring $R$ and $n>1$, the matrix ring $\mathbb{M}_{n}(R)$ has a maximal subring.
Finally, we prove some needed facts as follows. We remind the reader that if $T$ is a ring with $|T|<\mathfrak{a}$, where $\mathfrak{a}$ is an infinite cardinal number, then for each finitely generated left/right $T$-module $M$, we have $|M|<\mathfrak{a}$. To see this, note that since $M$ is finitely generated as $T$-module, we infer that there exists a natural number $n$ such that $M$ is a homomorphic image of $T^{n}$ (as left/right $T$-module). Hence $|M|\leq|T|^{n}<\mathfrak{a}^{n}=\mathfrak{a}$. Now the following is in order.
@@
Let $T$ be a left Noetherian ring and $I$ be a nilpotent ideal of $T$. If $\mathfrak{a}$ be an infinite cardinal number, then $|T|<\mathfrak{a}$ if and only if $|T/I|<\mathfrak{a}$.
@@
It is clear that, if $|T|<\mathfrak{a}$, then $|T/I|<\mathfrak{a}$ for each ideal $I$ of $T$. Conversely, assume that $I$ be a nilpotent ideal of $T$ and $|T/I|<\mathfrak{a}$. Since $I$ is a nilpotent ideal of $T$, we infer that there exists a natural number $n$ such that $I^{m}=0$. Now consider the chain $0=I^{m}\subseteq I^{m-1}\subseteq\cdots\subseteq I\subseteq T$. Since $T$ is a left Noetherian ring, we conclude that for each $i$, $I^{i}$ is a finitely generated left ideal. Hence for each $i$, the left $T$-module $I^{i}/I^{i+1}$, $0\leq i\leq m-1$ is finitely generated. Thus $|I^{i}/I^{i+1}|<\mathfrak{a}$. In particular, $|I^{m-1}|<\mathfrak{a}$. Therefore for each $i$ we deduce that $|I^{i}|<\mathfrak{a}$ and thus $|T|<\mathfrak{a}$. ∎
Note that in the proof of the previous lemma we only use the fact that $I$ is a nilpotent ideal of $T$ such that $I^{i}$ is a finitely generated left ideal of $T$, for each $i$. It is clear that by the condition of the previous lemma we conclude that $T$ is finite (resp. countable) if and only if $T/I$ is finite (resp. countable).
We conclude this section by a result about the structure of maximal subrings of a certain direct product of rings. If $A$ and $B$ be rings, then we say that $A$ and $B$ are homomorphically non-isomorphic rings, whenever for each proper ideals $I$ of $A$ and $J$ of $B$, the rings $A/I$ and $B/J$ are not isomorphic as ring. For example if $A=\mathbb{M}_{n_{1}}(R_{1})\times\cdots\times\mathbb{M}_{n_{k}}(R_{k})$, where $k\geq 1$, each $n_{i}\geq 2$ and $R_{i}$ are arbitrary rings and $B$ is any commutative ring, then $A$ and $B$ are homomorphically non-isomorphic rings (note that clearly, each homomorphic image of $A$ is a non-commutative ring). Now the following is in order.

### v2

@@ 2. Preliminaries
Let $T$ be a ring and $I$ and $J$ be left ideals of $T$. We say $I$ and $J$ are similar left ideals if $T/I\cong T/J$ as left $T$-modules. If $a\in T$, define $(I:a):=\{x\in T\ |\ xa\in I\}$, then it is not hard to see that $I$ and $J$ are similar if and only if there exists $c\in T$ such that $I+Tc=T$ and $J=(I:c)$, see [13, Proposition 1.3.6]. For a left ideal $I$ of $T$, we have the ring isomorphism $\mathbb{I}(I)/I\cong(\operatorname{End}_{T}(T/I))^{\text{op}}$, see [23, P. 14, 1.11]. Consequently, if $I$ and $J$ are similar left ideals of $T$, then $\mathbb{I}(I)/I\cong\mathbb{I}(J)/J$ as rings. Thus, up to ring isomorphism, the ring $\mathbb{I}(I)/I$ is uniquely determined by similarity classes; in particular $|\mathbb{I}(I)/I|=|\mathbb{I}(J)/J|$. In other words, the cardinality of $\mathbb{I}(I)/I$ is invariant under similarity. Now, let $M$ be a maximal left ideal of $T$. The following facts are useful:
@@
    If $N$ is a (maximal) left ideal of $T$ similar to $M$, then $N$ is an ideal of $T$ if and only if $N=M$, which in turn holds if and only if $M$ is an ideal of $T$.
To see the previous facts, note that $(1)$ follows directly from [13, Proposition 1.3.6], because obviously $M+Tc=T$. For $(2)$, assume that $N$ and $M$ are similar left ideals and that $N$ is an ideal. Since $T/N\cong T/M$ as left $T$-modules, we have $N=\operatorname{l.ann}_{T}(T/N)=\operatorname{l.ann}_{T}(T/M)\subseteq M$. Because $N$ is a maximal left ideal, the inclusion $N\subseteq M$ forces $N=M$. Hence $(2)$ holds. Consequently, if $M$ is a maximal left ideal of $T$ that is not an ideal of $T$, and if $[M]$ denotes the set of all (maximal) left ideals of $T$ similar to $M$, then by $(2)$ we have $M]\subseteq\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$, i.e., every (maximal) left ideal in $[M]$ fail to be a two-ideal. Clearly, similarity is an equivalence relation on the set of all left ideals of $T$, and in particular on $\operatorname{{Max}_{\ell}}(T)$. From $(1)$ and $(2)$, it is evident that for a maximal left ideal $M$, we have $|[M]|=1$ (i.e., $M]=\{M\}$) if and only if $M$ is an ideal of $T$ (and therefore in this case $M\in\operatorname{Max}(T)\cap\operatorname{{Max}_{r}}(T)$).
It is clear that for any left ideal $I$ of a ring $T$ and any $c\in T$, we have $(I:c)=T$ if and only if $c\in I$. Moreover, $I\subseteq(I:c)$ holds if and only if $c\in\mathbb{I}(I)$. In particular, if $M$ is a maximal left ideal of $T$, these observations immediately imply that $(M:c)=M$ if and only if $c\in\mathbb{I}(M)\setminus M$. Consequently, $M]=\{M\}\cup\{(M:c)\ |\ c\in T\setminus\mathbb{I}(M)\}$. Before starting our main result (which present a lower bound for $|[M]|$), we recall that, as noted earlier, for any $N\in[M]$, we have $|\mathbb{I}(M)/M|=|\mathbb{I}(N)/N|$. We now present the following key theorem.
@@
Assume that $T$ is a ring and $M\in\operatorname{{Max}_{\ell}}(T)$ is not an ideal of $T$ (i.e., $M$ is not a maximal ideal of $T$). Then $|[M]|\geq|\mathbb{I}(M)/M|+1$. In particular, $|\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)|\geq|\mathbb{I}(M)/M|+1$.
@@
Let $R=\mathbb{I}(M)$ and $D=R/M$. Since $M$ is not an ideal of $T$ we infer that $R$ is a proper subring of $T$. Take a fixed element $x\in T\setminus R$. We claim that the function $c+M\longmapsto(M:x+c)$ is a one-one function from $D$ to $M]\setminus\{M\}$. Since $x\in T\setminus R$ and $c\in R$ we deduce that $x+c\in T\setminus R$ and therefore $(M:x+c)\in[M]\setminus\{M\}$, by the previous observation. Hence it remain to show that the function is one-one. Hence assume that $a,b\in R$ and $(M:x+a)=(M:x+b)$, we must show that $a+M=b+M$, i.e., $a-b\in M$. Let $t\in(M:x+a)=(M:x+b)$, then we conclude that $tx+ta,tx+tb\in M$. It follows that $t(a-b)\in M$, i.e., $(M:x+a)\subseteq(M:a-b)$. Now if $a-b\notin M$, then we obtain that $a-b\in R\setminus M$, a fortiori (M:ab)=M(M:a-b)=M. This implies that $(M:x+a)\subseteq M$ and consequently by maximality of $(M:x+a)$ we deduce that $(M:x+a)=M$ which is absurd, because $x+a\in T\setminus R$. Thus $a-b\in M$, as desired. ∎
An immediate consequence of the previous theorem is that if $T$ is a ring and $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)\neq\emptyset$, then $|\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)|\geq 3$. In other words, if a ring $T$ possesses a maximal left ideal that is not a two-ideal, then $T$ has at least three distinct maximal left ideals that are not two-ideals.
@@
Let $T$ be a $K$-algebra over a field $K$. Then either $T$ is a left quasi-duo ring or $T$ has at least $|K|+1$ maximal left ideals which are not ideals of $T$ (i.e., $|\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)|\geq|K|+1$).
@@
Assume that $T$ is not a left quasi-duo ring and $M$ is a maximal left ideal of $T$ which is not an ideal of $T$. Since $K$ is a field which is contained in the center of $T$, we conclude that $K\subseteq R:=\mathbb{I}(M)$ and clearly $K\cap M=0$. Thus $K$ embeds in $R/M$ and we are done by Theorem 2.1. ∎
@@
Let $T$ be a $K$-algebra over an infinite field $K$. Then either $T$ is a left quasi-duo ring or $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$ is infinite (in fact, $|\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)|\geq|K|$).
@@
Thus, if $T$ is a ring and $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$ is finite, then for each $M\in\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$, we deduce that $\mathbb{I}(M)/M$ is a finite field. Moreover, if additionally, $T$ is a $K$-algebra over a field $K$ and $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$ is finite, then $K$ must be a finite field (note that for any maximal left ideal $M$ of $T$, the field $\mathbb{I}(M)/M$ contains a copy of $K$).
@@
Let $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)=\bigcup_{i\in I}[M_{i}]$ be the partition of $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$ into similarity classes. By Theorem 2.1, we obtain that $|\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)|\geq|I|+\sum_{i\in I}|\mathbb{I}(M)/M|$.
@@
It is clear from the above results that if $K$ is a field and $n>1$, then $|\operatorname{{Max}_{\ell}}(\mathbb{M}_{n}(K))|\geq|K|+1$ and $|\operatorname{{Max}_{r}}(\mathbb{M}_{n}(K))|\geq|K|+1$ (note, $\mathbb{M}_{n}(K)$ is a simple $K$-algebra and therefore each maximal left/right ideal of $\mathbb{M}_{n}(K)$ is not an ideal of $\mathbb{M}_{n}(K)$). In particular, if $K$ is an infinite field, then $\mathbb{M}_{n}(K)$ has at least $|K|$-many maximal left/right ideals (also see Remark 2.5). Next we employ the notation and results of [27] to show that if $R$ is an infinite division ring and $n>1$, then $T=\mathbb{M}_{n}(R)$ possesses infinitely many maximal left/right ideals. Recall that if $M$ is a maximal left ideal of a ring $R$ and $T=\mathbb{M}_{n}(R)$, then for each $u\in R^{n}\setminus M^{n}$, the set $D(M,u):=\{X\in T\ |\ Xu\in M^{n}\}$ is a maximal left ideal of $T$; indeed, $T/D(M,u)\cong(R/M)^{n}$ as left $T$-modules. Moreover, $\operatorname{{Max}_{\ell}}(T)=\{D(M,u)\ |\ M\in\operatorname{{Max}_{\ell}}(R),\ u\in R^{n}\setminus M^{n}\}$, see [27, Theorem 1.2]. In particular, when $R$ is a division ring, we obtain that $\operatorname{{Max}_{\ell}}(\mathbb{M}_{n}(R))=\{D(0,u)\ |\ 0\neq u\in R^{n}\}$.
@@
Hence to show that $\operatorname{{Max}_{\ell}}(\mathbb{M}_{n}(R))$ is infinite for an infinite division ring $R$, it suffices to produce infinitely many nonzero vectors $u\in R^{n}$ for which $D(0,u)$ are distinct. By [27, Proposition 2.3], if $M$ is a maximal left ideal of a ring $R$, $S=\mathbb{I}(M)$ its idealizer of $M$ in $R$, and $u,v\in S^{n}$, then $D(M,u)=D(M,v)$ (in $T=\mathbb{M}_{n}(R)$) if and only if there exists $c\in S\setminus M$ such that $v\equiv uc\ \mod(M)$, i.e., $v-uc\in M^{n}$. In particular, for a division ring $R$ the only possible maximal left ideals is $M=0$, so that $S=R$; consequently, for $u,v\in R^{n}$, we have $D(0,u)=D(0,v)$ if and only if there exists $0\neq c\in R$ such that $v=uc$. In other words $v$ and $u$ are right-parallel vectors in the right $R$-module $R^{n}$. Clearly, if $R$ is an infinite division ring and $n>1$, then as right $R$-module $R^{n}$ contains infinitely many right non-parallel vectors. This observations yield the following quick corollary.
@@
Let $R$ be an infinite division ring and $n>1$. Then $T=\mathbb{M}_{n}(R)$ has infinitely many maximal left (resp. right) ideals.
@@
Let $R$ be a ring, $n>1$, and $T=\mathbb{M}_{n}(R)$. It is obvious that $\operatorname{{Max}_{\ell}}(T)\cap\operatorname{Max}(T)=\emptyset=\operatorname{{Max}_{r}}(T)\cap\operatorname{Max}(T)$. Moreover, the map $M\longmapsto M^{t}$ (where $M^{t}:=\{m^{t}\ |\ m\in M\}$ is the transpose of $M$), is a one-to-one correspondence from $\operatorname{{Max}_{\ell}}(T)$ onto $\operatorname{{Max}_{r}}(T)$. In particular, $|\operatorname{{Max}_{\ell}}(T)|=|\operatorname{{Max}_{r}}(T)|$.
@@
For any ring $T$, we have the following immediate facts:
@@
    $\operatorname{{Max}_{\ell}}(T)\cap\operatorname{Max}(T)=\operatorname{{Max}_{r}}(T)\cap\operatorname{Max}(T)=\operatorname{{Prm}_{\ell}}(T)\cap\operatorname{{Max}_{\ell}}(T)=\operatorname{{Prm}_{r}}(T)\cap\operatorname{{Max}_{r}}(T)=\operatorname{{Prm}_{\ell}}(T)\cap\operatorname{{Max}_{r}}(T)=\operatorname{{Prm}_{r}}(T)\cap\operatorname{{Max}_{\ell}}(T)$.
@@
    $\operatorname{Max}(T)\setminus\operatorname{{Max}_{\ell}}(T)=\operatorname{Max}(T)\setminus\operatorname{{Max}_{r}}(T)$.
Next, let we recall some basic facts from maximal subrings, as well as several needed results from [1]. A proper subring $S$ of a ring $T$ is maximal if and only if for every $x\in T\setminus S$, one has $S[x]=T$. Moreover, if $R$ is a proper subring of $T$ and there exists $\alpha\in T$ such that $R[\alpha]=T$, then by a standard application of Zorn’s Lemma, $T$ possesses a maximal subring $S$ with $R\subseteq S$ and $\alpha\notin S$. In particular, if $R\subsetneq T$ is a ring extension which is finite as ring (or as module) over $R$, then $T$ has a maximal subring containing $R$. Now let $R$ be a ring and $\Delta:=\{(r,r)\ |\ r\in R\}$, be the diagonal subring of $T:=R\times R$. It is easy to see that $\Delta[(1,0)]=T$. Consequently, $T$ has a maximal subring containing $\Delta$. In fact, every subring $S$ of $T$ that contains $\Delta$ has the form $\Delta(I):=\Delta+I\times I$, where $I$ is an ideal of $R$. Moreover, $\Delta(I)$ is a maximal subring of $T$ if and only if $I$ is a maximal ideal of $R$, see [1, Theorem 2.7]. Therefore a ring $R$ is simple exactly when $\Delta$ is a maximal subring of $R\times R$. The following results from [1] will be useful.
@@
    [1, Theorem 4.1]. Let $T$ be a ring and $A$ be a maximal right/left ideal of $T$ that is not an ideal of $T$. Then the idealizer of $A$ is a maximal subring of $T$. In particular, every ring either has a maximal subring or is a quasi-duo ring.
@@
    [1, Proposition 4.2]. Let $R$ be a maximal subring of a ring $T$ that contains a maximal left/right ideal $A$ of $T$ that is not two-sided ideal of $T$. Then $R$ is precisely the idealizer of $A$ in $T$.
@@
    [1, Theorem 4.4]. Let $T$ be a ring that is not a division ring. If $T$ is left/right primitive ring (in particular, if $T$ is simple), then $T$ has a maximal subring.
@@
    [1, Theorem 3.1]. For every ring $R$ and every $n>1$, the matrix ring $\mathbb{M}_{n}(R)$ has a maximal subring.
Finally, we prove several auxiliary facts. Recall that if $T$ is a ring with $|T|<\mathfrak{a}$, where $\mathfrak{a}$ is an infinite cardinal number, then for each finitely generated left/right $T$-module $M$, we have $|M|<\mathfrak{a}$. Indeed, if $M$ is finitely generated, there exists a natural number $n$ such that $M$ is a homomorphic image of $T^{n}$ (as left/right $T$-module); hence $|M|\leq|T|^{n}<\mathfrak{a}^{n}=\mathfrak{a}$.
@@
Let $T$ be a left Noetherian ring and $I$ a nilpotent ideal of $T$. If $\mathfrak{a}$ is an infinite cardinal number, then $|T|<\mathfrak{a}$ if and only if $|T/I|<\mathfrak{a}$.
@@
The implication $|T|<\mathfrak{a}\ \Longrightarrow|T/I|<\mathfrak{a}$ is obvious for every ideal $I$ of $T$. Conversely, assume $I$ is a nilpotent ideal of $T$ and $|T/I|<\mathfrak{a}$. Since $I$ is nilpotent, there exists a natural number $n$ with $I^{m}=0$. Consider the chain $0=I^{m}\subseteq I^{m-1}\subseteq\cdots\subseteq I\subseteq T$. Because $T$ is left Noetherian, each $I^{i}$ is a finitely generated left ideal. Consequently, for each $i$, $0\leq i\leq m-1$, the left $T$-module $I^{i}/I^{i+1}$ is finitely generated, and therefore $|I^{i}/I^{i+1}|<\mathfrak{a}$. In particular, $|I^{m-1}|<\mathfrak{a}$. By descending induction we obtain $|I^{i}|<\mathfrak{a}$, for all $i$, and finally $|T|<\mathfrak{a}$. ∎
Note that in the proof of the previous lemma we only used that $I$ is a nilpotent ideal whose powers $I^{i}$ are finitely generated left ideals of $T$. Under the hypotheses of the lemma, we immediately conclude that $T$ is finite (resp. countable) precisely when $T/I$ is finite (resp. countable).
We close this section with a result concerning the structure of maximal subrings of certain direct products of rings. Two rings $A$ and $B$ are called homomorphically non-isomorphic, if for every proper ideal $I$ of $A$ and every proper ideal $J$ of $B$, the rings $A/I$ and $B/J$ are not isomorphic. For example, if $A=\mathbb{M}_{n_{1}}(R_{1})\times\cdots\times\mathbb{M}_{n_{k}}(R_{k})$, where $k\geq 1$, each $n_{i}\geq 2$ and $R_{i}$ are arbitrary rings and $B$ is any commutative ring, then $A$ and $B$ are homomorphically non-isomorphic (since every homomorphic image of $A$ is noncommutative). With this preparation we state the following.